7. Suppose sequence (-1 +1 –3 +1 –1 –3 +1 +1) is transmitted and the chip sequences for four users are: a. –1 -1 -1 +1 +1 -1 +1 +1 b. –1 -1 +1 -1 +1 +1 +1 -1 c. -1 +1 -1 +1 +1 +1 -1 -1 d. -1 +1 -1 -1 -1 -1 +1 -1 Determine the signal that each user sent (assume that each user can transmit – 1, 0, and +1)
Answer #1

Solution

7)

To determine the signal that each user sent, we can compute each of the
normalized inner products

a. (-1 +1 –3 +1 –1 –3 +1 +1) * (-1 -1 -1 +1 +1 -1 +1 +1)/8 = 8/8 = +1
b. (-1 +1 –3 +1 –1 –3 +1 +1) * (-1 -1 +1 -1 +1 +1 +1 -1)/8 = -8/8 = -1
c. (-1 +1 –3 +1 –1 –3 +1 +1) * (-1 +1 -1 +1 +1 +1 -1 -1)/8 = 0/8 = 0
d. (-1 +1 –3 +1 –1 –3 +1 +1) * (-1 +1 -1 -1 -1 -1 +1 -1)/8 = 8/8 = +1

---

all the best

please upvote